# arXiv:2603.01198v1[eess.SY] 01 Mar 2026

### v1

@@ 3.2 Baseline operating conditions
Figure 3 presents the annual operating profile of the Frontier hot-temperature water (HTW) cooling loop across 2023. Three key observations emerge from the data. First, the computational heat load remains relatively stable between 7 and 12 MW throughout the year (Figure 3(a)), reflecting the sustained utilization typical of leadership-class supercomputers. Second, the supply temperature $T_{\mathrm{supply}}$ varies from approximately 12C in winter to 29C in summer, tracking the ambient wet-bulb temperature with a roughly constant heat exchanger approach (Figure 3(b)). The return temperature $T_{\mathrm{return}}$ follows a similar seasonal pattern, with 7-day averages ranging from 27C to 35C. The dashed horizontal line at 42C indicates the optimization constraint threshold, set 3C below the 45C equipment limit. Third, and most notably, the pump flow rate exhibits a pronounced step change around May (Figure 3(c)), jumping from approximately 200 kg/s to 350–400 kg/s and remaining elevated through October. This step-like behavior, rather than a smooth modulation proportional to cooling demand, suggests the pumps are operated at fixed-speed setpoints without continuous optimization.
@@
3.3 Over-pumping diagnosis
To quantify the extent of excess pumping, we compare the measured baseline flow rate against the analytical minimum flow required to satisfy the thermal constraint $T_{\mathrm{return}}\leq 42^{\circ}$C at each CDU branch. The minimum safe flow is computed from the energy balance as:
$\dot{m}_{\mathrm{min}}=\max_{j\in\{1,2,3\}}\frac{Q_{j}}{f_{j}\,c_{p}\,(T_{\mathrm{limit}}-T_{\mathrm{supply}})}$ (18)
where $Q_{j}$ and $f_{j}$ denote the heat load and flow fraction for CDU branch $j$, respectively.
Figure 4 presents the results of this comparison. Figure 4(a) plots the measured baseline flow against the computed minimum safe flow for all 47,186 operating points. Nearly all points lie above the 1:1 line, confirming systematic over-pumping. The dashed $1.5\times$ line captures the central tendency of the data. Figure 4(b) shows the histogram of the over-pumping ratio (baseline divided by minimum), with a median of $1.5\times$. This indicates that, on average, the facility pumps 50% more coolant than thermally necessary. Because pump power scales as $P\propto\dot{m}^{n}$ with $n\in[2,3]$ under the affinity laws, even a modest flow reduction yields substantial energy savings: a $1.5\times$ flow reduction corresponds to a $1.5^{3}=3.4\times$ reduction in pump power under the cubic affinity law.
@@ 3.4 Annual energy comparison across strategies
Four control policies are compared: (1) Baseline (current operations), (2) Strategy A (pump flow optimization only), (3) Strategy B (joint flow and supply temperature setpoint co-optimization with unconstrained ramp rates), and (4) Strategy C (ramp-constrained co-optimization, applying the same objective as Strategy B but enforcing ramp-rate limits of $\pm$50 kg/s per step for flow and $\pm$1C per step for $T_{\mathrm{supply}}$). Strategy C follows model predictive control (MPC) principles within the digital twin framework, using the calibrated $\varepsilon$-NTU model as the prediction engine while imposing operationally realistic actuator constraints. Table 3 summarizes the annual energy results and Figure 5 provides a visual comparison.
@@
A central finding is that cooling tower (CT) fan energy dominates the total cooling energy budget, accounting for 73% of baseline consumption (1,325 MWh out of 1,814 MWh). This 2.7:1 ratio between CT fan and pump energy has important strategic implications. Strategy A, which optimizes only pump flow, achieves a 75.7% reduction in pump energy but leaves the larger CT component untouched, yielding 20.4% total savings. Strategy B introduces supply temperature setpoint reset as an additional decision variable, enabling the optimizer to raise $T_{\mathrm{supply}}$ by an average of 2.7C. This reduces the CT fan workload by 17.9% while accepting slightly higher pump energy than Strategy A (63.1% vs. 75.7% pump savings). The net result is a 30.1% total savings for Strategy B, representing a 48% improvement over the flow-only approach. Strategy C, which enforces operationally realistic ramp-rate constraints, achieves 27.8% total savings, recovering 92.4% of the unconstrained theoretical bound established by Strategy B.
3.5 Monthly and seasonal energy profiles
Figure 6 presents the monthly total cooling energy for all four strategies. The savings are not uniformly distributed across the year. Winter months (January through March) exhibit savings of approximately 15–20%, constrained by the already-low supply temperatures that limit room for setpoint adjustment. Summer months (May through September) realize the largest absolute savings, where the optimizer simultaneously reduces excess flow and raises $T_{\mathrm{supply}}$ to avoid unnecessary subcooling by the cooling tower. The peak monthly saving of 33% occurs in July, when the baseline cooling load is highest.
@@ 3.5 Monthly and seasonal energy profiles
Figure 7 provides a compact overview of the savings across all months and strategies. The color-coded heatmap reveals a clear seasonal structure: Strategy A contributes minimally during winter (6–10%) but achieves 25–37% during summer, when the baseline flow is most inflated relative to thermal need. Strategies B and C consistently outperform Strategy A across all months, with the largest absolute advantage during the shoulder seasons (March through April, October through November) where setpoint reset remains effective even when pump savings are modest. Notably, the gap between Strategy B and Strategy C is smallest during summer (1–3%), indicating that the ramp-rate constraints impose the least penalty precisely when savings are highest.
Figure 7: Monthly total cooling energy savings (%) by control policy. Color intensity indicates magnitude. Co-optimization strategies (B and C) outperform flow-only optimization (A) across all months, with the advantage most pronounced during shoulder and winter seasons.
Table 4 quantifies the seasonal breakdown. The summer period (June through August) accounts for 43% of total annual savings under Strategy C despite spanning only 25% of the calendar year. The implementability cost of ramp-rate constraints, defined as the gap between Strategies B and C, varies from 1.5% in winter to 2.6% in summer. The relative cost is largest during winter, when smaller absolute savings amplify the impact of constraint-induced lag.
@@
3.6 Supply temperature setpoint mechanism
The mechanism by which supply temperature co-optimization achieves additional savings is illustrated in Figure 8. Figure 8(a) shows the distribution of $\Delta T_{\mathrm{supply}}$ (optimized minus baseline) across all operating points. The optimizer raises $T_{\mathrm{supply}}$ by a mean of $2.7^{\circ}$C in 97.9% of timesteps, while lowering it in only 2.0% of cases (typically during transient high-load events). Figure 8(b) compares the monthly mean $T_{\mathrm{supply}}$ for Baseline, Strategy B, and Strategy C. The optimized strategies consistently maintain a higher setpoint than the baseline, with the gap widening in summer when the trade-off between pump and CT energy is most consequential. During winter, the baseline supply temperature is already near the cooling water temperature, leaving limited room for further adjustment; Strategies B and C converge near the baseline values.
The physical interpretation is that the existing control system overcools the supply water below what is thermally necessary, forcing the cooling tower to work harder to produce unnecessarily cold cooling water. By raising $T_{\mathrm{supply}}$, the optimizer accepts a warmer supply stream, which reduces the CT approach temperature and corresponding fan power, at the cost of requiring slightly more pump flow to maintain the return temperature below 42C. The net energy trade-off favors the raised setpoint in nearly all conditions.
@@ 3.6 Supply temperature setpoint mechanism
3.7 Operating envelope shift
Figure 9 presents the joint distribution of flow rate and supply temperature for Baseline, Strategy B, and Strategy C, visualized through kernel density contours overlaid on the raw operating points. Under baseline operations (Figure 9(a)), the facility occupies a broad operating envelope centered at 257 kg/s mean flow and 20.1C mean supply temperature, with a characteristic bimodal structure reflecting the seasonal flow step change observed in Figure 3.
Strategy B (Figure 9(b)) shifts the operating envelope toward lower flow rates ($\mu=187$ kg/s) and higher supply temperatures ($\mu=22.8^{\circ}$C), compressing the flow distribution and eliminating the high-flow summer mode. Strategy C (Figure 9(c)) exhibits a nearly identical centroid ($\mu=189$ kg/s, $22.6^{\circ}$C) but with a slightly broader distribution, reflecting the smoothing effect of ramp-rate constraints that prevent the optimizer from reaching the instantaneous setpoint at every timestep. The similarity between the Strategy B and Strategy C envelopes confirms that the ramp constraints primarily affect transition dynamics rather than the steady-state operating target.
Figure 9: Joint distribution of flow rate and supply temperature for (a) Baseline, (b) Strategy B, and (c) Strategy C. Contour lines show kernel density estimates; dotted lines indicate mean values. Co-optimization shifts the operating envelope toward lower flow and higher supply temperature, while Strategy C closely tracks the Strategy B centroid with modestly broader spread due to ramp-rate constraints.
@@ 3.8 Strategy C dynamic behavior under constraints
Figure 10 presents a representative summer week (168 hours in July), comparing the dynamic trajectories of Baseline, Strategy B, and Strategy C across four variables. Figure 10(a) shows the flow rate: the baseline operates near 400 kg/s with abrupt step changes, while both optimized strategies modulate flow between 150 and 400 kg/s in response to load variations. The Strategy C trajectory is smoother than Strategy B, as the ramp-rate constraints prevent large instantaneous jumps. Figure 10(b) shows $T_{\mathrm{supply}}$: Strategy C maintains a consistently higher setpoint (28–32C) compared to the baseline (24–28C), reflecting the setpoint reset mechanism. Figure 10(c) demonstrates the resulting total power reduction, with Strategy C achieving 30–50% lower instantaneous power during most of the week. Figure 10(d) confirms that Strategy C operates near the 42C return temperature constraint throughout, with the few exceedances occurring exclusively at the maximum pump capacity (420 kg/s), where the system has exhausted its flow control authority. All operating points remain below the 45C equipment limit.
@@
3.9 Ramp-rate compliance
A key practical concern for real-world deployment is whether the optimization commands can be tracked by physical actuators without inducing water hammer, thermal shock, or control instability. Figure 11 directly addresses this by plotting the absolute step-to-step change in flow rate (Figure 11(a)) and $T_{\mathrm{supply}}$ setpoint (Figure 11(b)) for both the unconstrained Strategy B and the ramp-constrained Strategy C.
Strategy B produces flow rate changes up to 189 kg/s per 10-minute step and temperature changes up to 4.4C per step, frequently violating both the $\pm$50 kg/s and $\pm$1C limits that represent conservative actuator and process constraints. These ramp-rate limits are informed by typical variable-frequency drive slew rates and thermal inertia considerations for large-scale chilled and hot water distribution systems (ASHRAE, 2020). Strategy C, by contrast, respects both ramp-rate limits by construction. Temperature setpoint changes under Strategy C are strictly bounded below $\pm$1C, and flow rate changes remain within the $\pm$50 kg/s envelope throughout normal operation. This constraint satisfaction comes at a modest cost: Strategy C achieves 27.8% total savings versus 30.1% for Strategy B, retaining 92.4% of the unconstrained theoretical bound.
Figure 11: Step-to-step ramp-rate comparison between unconstrained Strategy B and ramp-constrained Strategy C during a summer week: (a) flow rate changes with $\pm$50 kg/s limit, and (b) $T_{\mathrm{supply}}$ setpoint changes with $\pm$1C limit. Strategy C respects both operational constraints while Strategy B frequently violates them.
3.10 Pump power duration curves
Figure 12 presents the annual pump power duration curves for all four strategies, providing an alternative perspective on the savings distribution across operating hours. The baseline curve shows pump power exceeding 100 kW for approximately 55% of the year, with a peak near 200 kW. All three optimized strategies shift the curve downward substantially. Strategy A (flow only) achieves the largest pump power reduction, since it minimizes flow without regard for CT energy. Strategies B and C maintain slightly higher pump power levels (as they trade pump energy for CT savings), but the overall shift relative to the baseline remains large. Notably, the Strategy C curve lies slightly above Strategy B in the 20–60% exceedance range, reflecting the ramp-rate constraints that prevent Strategy C from reaching the instantaneous optimum at every timestep.
Figure 12: Annual pump power duration curves showing the distribution of pump power across all operating hours. All optimization strategies substantially reduce pump power relative to the baseline, with the greatest reductions during high-load summer operation.
@@ 3.11 Sensitivity to pump curve exponent
The pump power model $P_{\mathrm{pump}}=P_{\mathrm{pump,nom}}(\dot{m}/\dot{m}_{\mathrm{nom}})^{n}$ Equation (4) depends on the exponent $n$, which ranges from $n=2$ (conservative estimate with residual static head) to $n=3$ (standard affinity law for fully turbulent flow). Because the exact Frontier pump curves are not publicly available, we evaluate all strategies across three representative exponents. Table 5 and Figure 13 present these results.
@@
The sensitivity analysis reveals two important insights. First, the total savings increase monotonically with the pump exponent across all strategies, as expected from the nonlinear relationship between flow and power. Under the most conservative assumption ($n=2$), Strategy C still achieves 21.3% total savings; under the standard affinity law ($n=3$), the savings reach 27.8%. This 21–28% range provides a robust bound on the expected benefits. Second, the relative advantage of co-optimization (Strategies B and C) over flow-only optimization (Strategy A) is most pronounced at lower pump exponents. At $n=2$, Strategy C nearly doubles the savings of Strategy A (21.3% vs. 10.9%), because when pump power is less sensitive to flow, the CT fan savings from setpoint reset become the dominant contributor. At $n=3$, Strategy C achieves a 36% relative improvement over Strategy A (27.8% vs. 20.4%). This demonstrates that supply temperature co-optimization is valuable across the full range of plausible pump characteristics.
Figure 13: Sensitivity of total cooling energy savings to pump curve exponent for all three optimization strategies. Co-optimization (Strategies B and C) consistently outperforms flow-only optimization across all exponents.
@@ 4.1 System-level perspective on cooling system optimization
A primary contribution of this work is the demonstration that optimizing pump flow alone captures only a fraction of the available cooling energy savings. The finding that CT fan energy constitutes 73% of the total cooling energy budget fundamentally shifts the optimization landscape. While the pump affinity law ($P\propto\dot{m}^{3}$) makes flow reduction an attractive target in isolation, the system-level optimum lies in coordinating pump flow and supply temperature setpoint to minimize the combined pump and CT fan consumption. This coordination produces a counterintuitive outcome: the co-optimized Strategy B uses more pump energy than the flow-only Strategy A (180,174 vs. 118,772 kWh), yet achieves greater total savings (30.1% vs. 20.4%) by reducing the dominant CT fan component.
This finding aligns with the broader principle that component-level optimization can be suboptimal when interdependencies exist between subsystems. In data center cooling, lowering the supply temperature requires the cooling tower to produce colder water, which demands more fan energy. The digital twin captures these interdependencies through the coupled $\varepsilon$-NTU heat exchanger model and the cooling tower performance curves, enabling the optimizer to navigate the pump-versus-CT trade-off at each operating point. The operating envelope analysis (Figure 9) further illustrates this rebalancing: the shift from the baseline centroid (257 kg/s, 20.1C) to the co-optimized centroid (187–189 kg/s, 22.6–22.8C) reflects a deliberate reallocation of energy between the two subsystems rather than simple minimization of either one.
@@ 4.2 Practical implementability of the ramp-constrained strategy
Strategy C addresses a gap often observed in optimization studies that report unconstrained theoretical savings without considering actuator limitations. The $\pm$50 kg/s flow rate limit and $\pm$1C setpoint change limit per 10-minute step represent conservative bounds informed by typical variable-frequency drive slew rates and thermal inertia considerations for large-scale cooling systems (ASHRAE, 2020). Strategy C recovers 92.4% of the unconstrained savings established by Strategy B (27.8% vs. 30.1%) while producing smooth, actuator-friendly control trajectories. The 2.3% reduction from the theoretical bound is the cost of implementability, a modest penalty for a strategy that can be deployed on real hardware without risking water hammer, thermal shock, or control oscillations.
The seasonal analysis (Table 4) provides additional nuance: the implementability gap ranges from 1.5% in winter to 2.6% in summer in absolute terms, though the relative impact is proportionally larger during winter when the total savings base is smaller. This pattern arises because winter conditions offer limited $T_{\mathrm{supply}}$ headroom, and the ramp constraints slow convergence to the optimal setpoint during brief warm-weather excursions. A seasonally adaptive ramp-rate policy could further narrow this gap, though such refinement is left for future work.
The nine thermal exceedances observed under Strategy C (Table 3) all occur at the maximum pump capacity of 420 kg/s during extreme summer load peaks, where the system has exhausted its flow control authority. The maximum return temperature of 42.9C remains well below the 45C equipment limit, and the exceedances represent 0.02% of all operating points. These capacity-limited events are inherent to the system design rather than a deficiency of the optimization strategy, and the 3C safety margin between the optimization constraint (42C) and the equipment limit (45C) provides adequate defense-in-depth against such excursions.
4.3 Robustness across pump curve uncertainty
The 21–28% savings range across pump exponents (Table 5) provides confidence that the proposed approach is robust to uncertainty in the pump power model. This is practically significant because detailed pump curves for installed equipment are often unavailable to researchers and operators. Even under the most conservative assumption ($P\propto\dot{m}^{2}$), Strategy C achieves over 21% total cooling energy savings, representing approximately 380 MWh of annual reduction for the Frontier facility. Under the typical assumption ($P\propto\dot{m}^{2.5}$), the savings reach 24.2%, corresponding to approximately 435 MWh annually. At utility rates typical for the Oak Ridge area, these savings translate to an estimated annual cost reduction on the order of \$30,000–\$50,000, with the additional benefit of reduced carbon emissions proportional to the local grid emission factor.
4.4 Scalability and generalizability
The methodology is applicable to other liquid-cooled data centers that employ external heat exchangers and cooling towers, provided that sufficient operational data are available for digital twin calibration. The Modelica-based modeling framework enables component substitution (e.g., replacing the counterflow heat exchanger with a crossflow configuration) without modifying the optimization algorithm. The $\varepsilon$-NTU analytical model used within the Strategy C optimizer is computationally lightweight, solving each 10-minute optimization in under 10 milliseconds on a standard workstation. This computational efficiency makes the approach suitable for deployment as an online supervisory controller with a 10-minute update cycle.

### v2

@@ 3.2 Baseline operating conditions
Figure 3 presents the annual operating profile of the Frontier HTW cooling loop across 2023. The computational heat load remains relatively stable between 7 and 12 MW year-round, reflecting the sustained utilization typical of leadership-class supercomputers. The supply temperature tracks the ambient wet-bulb temperature seasonally (12C in winter to 29C in summer), and the return temperature remains well below the 42C optimization constraint at all times. Most notably, the pump flow rate exhibits a pronounced step change around May, jumping from approximately 200 kg/s to 350–400 kg/s and remaining elevated through October. This step-like behavior suggests the pumps are operated at fixed-speed setpoints without continuous optimization.
3.3 Over-pumping diagnosis
To quantify excess pumping, the measured baseline flow is compared against the analytical minimum flow required to satisfy $T_{\mathrm{return}}\leq 42$ C at each CDU branch:
$\dot{m}_{\mathrm{min}}=\max_{j\in\{1,2,3\}}\frac{Q_{j}}{f_{j}\,c_{p}\left(T_{\mathrm{limit}}-T_{\mathrm{supply}}\right)}$ (18)
[…]
@@ 3.4 Annual energy comparison across strategies
Table 2 summarizes the annual energy results for all four strategies and Figure 5 provides a visual comparison. A central finding is that cooling tower (CT) fan energy dominates the total cooling energy budget, accounting for 73% of baseline consumption (1,325 MWh out of 1,814 MWh). Strategy A optimizes only pump flow, achieving a 75.7% reduction in pump energy but leaving the CT component untouched (20.4% total savings). Strategy B introduces supply temperature setpoint reset, raising $T_{\mathrm{supply}}$ by an average of 2.7C, which reduces CT fan energy by 17.9% and achieves 30.1% total savings. Strategy C enforces ramp-rate constraints and achieves 27.8% total savings, recovering 92.4% of the unconstrained theoretical bound.
3.5 Monthly energy profiles and supply temperature mechanism
Figure 6 presents the monthly total cooling energy for all four strategies. Winter months (January through March) exhibit savings of 15–20%, constrained by already-low supply temperatures that limit setpoint adjustment. Summer months (May through September) realize the largest savings, with a peak of 33% in July. The monthly savings heatmap (Figure A1 in Appendix A) confirms that co-optimization strategies outperform flow-only optimization across all months, with the B-to-C gap smallest in summer (1–3%).
Table 3 quantifies the seasonal breakdown. Summer (June through August) accounts for 43% of total annual savings under Strategy C despite spanning only 25% of the calendar year. The implementability cost varies from 1.5% in winter to 2.6% in summer; the relative cost is largest in winter, where smaller absolute savings amplify the impact of constraint-induced lag.
The mechanism behind co-optimization savings is illustrated in Figure 7. The optimizer raises $T_{\mathrm{supply}}$ by a mean of +2.7+2.7 C in 97.9% of timesteps, with the gap widening in summer when the pump-versus-CT trade-off is most consequential. During winter, the baseline supply temperature is already near the cooling water temperature, leaving limited room for adjustment.
@@ 3.5 Monthly energy profiles and supply temperature mechanism
3.6 Operating envelope shift
Co-optimization shifts the operating envelope from the baseline centroid (257 kg/s, 20.1C) toward lower flow and higher supply temperature: Strategy B centers at 187 kg/s and 22.8C, while Strategy C centers at 189 kg/s and 22.6C. The joint distribution plots (Figure A2 in Appendix A) confirm that both strategies eliminate the baseline bimodal flow structure and compress the operating envelope. The near-identical centroids of Strategies B and C indicate that the ramp constraints primarily affect transition dynamics rather than the steady-state operating target.
@@ 3.7 Strategy C dynamic behavior under constraints
Figure 8 presents a representative summer week (168 hours in July), comparing the dynamic trajectories of Baseline, Strategy B, and Strategy C. Strategy C produces smoother control trajectories than Strategy B while maintaining a consistently higher supply temperature setpoint (28–32C vs. baseline 24–28C), achieving 30–50% lower instantaneous total power during most of the week. The return temperature operates near the 42C constraint throughout, with the few exceedances occurring exclusively at maximum pump capacity (420 kg/s) where the system has exhausted its flow control authority. All operating points remain below the 45C equipment limit.
@@
3.8 Ramp-rate compliance and pump power distribution
A practical concern for real-world deployment is whether optimization commands can be tracked by physical actuators. Strategy B produces flow rate changes up to 189 kg/s per step and temperature changes up to 4.4C per step, frequently violating the $\pm 50$ kg/s and $\pm 1$ C limits that represent conservative actuator constraints informed by variable-frequency drive slew rates (ASHRAE, 2020). Strategy C respects both limits by construction (Figure A3 in Appendix A), achieving 27.8% total savings versus 30.1% for Strategy B. The annual pump power duration curves (Figure A4 in Appendix A) confirm that all optimized strategies substantially reduce pump power relative to the baseline, with Strategy A achieving the largest pump reduction and Strategies B and C trading slightly higher pump power for cooling-tower fan savings.
@@ 3.9 Sensitivity to pump curve exponent
Because the exact Frontier pump curves are not publicly available, all strategies are evaluated across three representative exponents ($n\in\{2.0,2.5,3.0\}$). Table 4 presents the results, while Figure A5 in Appendix A provides a visual comparison. Under the most conservative assumption ($n=2$), Strategy C achieves 21.3% total savings; under the standard affinity law ($n=3$), savings reach 27.8%. This 21–28% range provides a robust bound on expected benefits. The relative advantage of co-optimization over flow-only optimization is most pronounced at lower pump exponents: at $n=2$, Strategy C nearly doubles the savings of Strategy A (21.3% vs. 10.9%), because cooling-tower fan savings from setpoint reset become the dominant contributor when pump power is less sensitive to flow.
@@ 4.1 System-level perspective on cooling system optimization
A primary contribution of this work is the demonstration that optimizing pump flow alone captures only a fraction of the available cooling energy savings. The finding that CT fan energy constitutes 73% of the total cooling energy budget fundamentally shifts the optimization landscape. While the pump affinity law ($P\propto\dot{m}^{3}$) makes flow reduction an attractive target in isolation, the system-level optimum lies in coordinating pump flow and supply temperature setpoint to minimize the combined pump and CT fan consumption. This coordination produces a counterintuitive outcome: the co-optimized Strategy B uses more pump energy than the flow-only Strategy A (180,174 vs. 118,772 kWh), yet achieves greater total savings (30.1% vs. 20.4%) by reducing the dominant CT fan component. This finding aligns with the broader principle that component-level optimization can be suboptimal when interdependencies exist between subsystems, a pattern likely to recur in other coupled thermal systems where multiple energy-consuming components share hydraulic and thermal coupling.
@@ 4.2 Practical implementability of the ramp-constrained strategy
Strategy C addresses a gap often observed in optimization studies that report unconstrained theoretical savings without considering actuator limitations. The $\pm 50$ kg/s flow rate limit and $\pm 1$ C setpoint change limit per 10-minute step represent conservative bounds informed by typical variable-frequency drive slew rates and thermal inertia considerations for large-scale cooling systems (ASHRAE, 2020). Strategy C recovers 92.4% of the unconstrained savings (27.8% vs. 30.1%) while producing smooth, actuator-friendly control trajectories. The 2.3% reduction from the theoretical bound is a modest penalty for a strategy deployable on real hardware without risking water hammer, thermal shock, or control oscillations. The seasonal analysis (Table 3) provides additional nuance: the implementability gap ranges from 1.5% in winter to 2.6% in summer, with the relative impact proportionally larger during winter when limited $T_{\mathrm{supply}}$ headroom slows convergence to the optimal setpoint. A seasonally adaptive ramp-rate policy could further narrow this gap.
4.3 Robustness and generalizability
The 21–28% savings range across pump exponents (Table 4) provides confidence that the proposed approach is robust to uncertainty in the pump power model, which is practically significant because detailed pump curves for installed equipment are often unavailable. Even under the most conservative assumption ($P\propto\dot{m}^{2}$), Strategy C achieves over 21% total cooling energy savings (approximately 380 MWh annually for the Frontier facility). The methodology is applicable to other liquid-cooled data centers that employ external heat exchangers and cooling towers, provided that sufficient operational data are available for digital twin calibration. The Modelica-based framework enables component substitution (e.g., replacing counterflow with crossflow heat exchangers) without modifying the optimization algorithm, and the $\varepsilon$-NTU analytical model solves each 10-minute optimization in under 10 ms, making the approach suitable for online supervisory control.